# Evaluating Transparency of Machine Generated Fact Checking Explanations

### v1

@@ Abstract
An important factor when it comes to generating fact-checking explanations is the selection of evidence: intuitively, high-quality explanations can only be generated given the right evidence. In this work, we investigate the impact of human-curated vs. machine-selected evidence for explanation generation using large language models. To assess the quality of explanations, we focus on transparency (whether an explanation cites sources properly) and utility (whether an explanation is helpful in clarifying a claim). Surprisingly, we found that large language models generate similar or higher quality explanations using machine-selected evidence, suggesting carefully curated evidence (by humans) may not be necessary. That said, even with the best model, the generated explanations are not always faithful to the sources, suggesting further room for improvement in explanation generation for fact-checking. Code and data are available here: https://github.com/ruixing76/Transparent-FCExp.
Evaluating Transparency of Machine Generated Fact Checking Explanations
@@ 1 Introduction
Automated fact-checking systems are designed to classify claims based on their truthfulness (Guo et al., 2022). However, debunking via simply calling it “false” can trigger a backfire effect, where the belief of false claims is further reinforced rather than hindered, reducing their effectiveness (Lewandowsky et al., 2012). As such, there’s growing research on generating textual explanations to justify the outcomes of fact-checking systems (Guo et al., 2022; Russo et al., 2023).
@@
Large language models (LLMs) have been used to generate explanations in various contexts (Wiegreffe et al., 2022). For automated fact-checking, the typical input is a claim and a list of retrieved evidence passages, from which a subset of evidence passages is selected and fed into the LLM for explanation generation. We present Figure 1 to illustrate the fact-checking explanation generation process.
How to perform effective evidence selection is an important question. Intuitively, high-quality explanations can only be generated given the right evidence. This constitutes the core research question of this paper: is carefully hand-curated evidence needed for explanation generation, or does machine-selected evidence suffice?
To answer this question, we need to first define how we evaluate the quality of explanations. We focus on two traits: transparency and utility. Transparency assesses whether an explanation cites the sources evidence appropriately Gao et al. (2023); this is motivated by journalistic practices in fact-checking, where experts carefully track their sources to ensure every cited statement is precise (Baker and Fairbank, 2022). Utility, on the other hand, captures whether users find an explanation helpful in clarifying the claim Liu et al. (2023). We evaluate transparency and utility manually. For transparency, we introduce a new approach where we mask citations in explanations and ask humans to recover them: the idea is that explanations that cite the sources correctly should produce a higher identification accuracy.
Our findings are: (1) machine-selected evidence is either on-par or superior to manually-selected evidence for explanation generation, depending on the LLM; (2) machine-selected evidence, based on manual analysis, appears to include a larger set of evidence passages that are relevant to the claim; and (3) despite encouraging performance using LLMs for generating fact-checking explanations, there is still room for improvement.
@@
Figure 1: Automated explanation generation for fact-checking. Given a claim and a list of evidence passages, a subset of these passages is selected, either by humans or machines, and input into a large language model (LLM) along with the claim to generate the explanation.
2 Automatic Explanation Generation
@@
Our LLM-based explanation generation system assumes the following as input: (1) a claim; (2) a (large) set of evidence passages, some of which may be useful for debunking the claim; and (3) the veracity of the claim (e.g. True). [^1] The rationale for assuming the veracity label as input is we see our explanation generation system being applied to a fact-checking classification system to provide a layer of interpretability in a real-world application. Because the set of evidence passages is typically large and not all of them are relevant, we propose a pipeline approach where we first do evidence selection and then feed the selected evidence, claim and veracity label to an LLM to generate the explanation. Figure 1 illustrates the whole process.
@@
We experiment with two methods: (1) (one-shot) prompt an LLM (Appendix Table 6); or (2) ask human judges to select a subset of the most relevant evidence passages given a claim and the full set of evidence passages. [^2] The dataset we use for our experiments contains the human-selected evidence passages, so we manually source this ourselves (see Section A).
@@
Given a claim, veracity label and selected evidence (machine- or human-selected), we (zero-shot) prompt an LLM to generate an explanation to clarify the claim (Appendix Table 5); the prompt explicitly asks it to use in-line citations, making it clear that all given evidence passages should be used (see “Generated Explanation” in Figure 1). We test a range of LLMs in our experiments, noting that we always use the same LLM for evidence selection and explanation generation.
Setting Model Precision Recall F1 Entropy Utility
Human GPT4 0.62$\pm$0.29 0.67$\pm$0.29 0.63$\pm$0.29 0.28$\pm$0.18 66.86$\pm$19.38
GPT35 0.52$\pm$0.29 0.59$\pm$0.30 0.52$\pm$0.29 0.41$\pm$0.18 72.47$\pm$19.52
LLaMA2-70b 0.48$\pm$0.31 0.52$\pm$0.32 0.49$\pm$0.31 0.34$\pm$0.16 65.13$\pm$18.00
Machine GPT4 0.72$\pm$0.32 0.79$\pm$0.31 0.74$\pm$0.31 0.18$\pm$0.19 76.34$\pm$17.88
GPT35 0.55$\pm$0.39 0.55$\pm$0.39 0.53$\pm$0.37 0.21$\pm$0.18 70.47$\pm$20.89
LLaMA2-70b 0.49$\pm$0.39 0.51$\pm$0.40 0.49$\pm$0.38 0.21$\pm$0.17 66.20$\pm$21.84
Table 1: Evaluation results of generated explanations for transparency (“Precision”, “Recall”, “F1” and “Entropy”) and utility (“Utility”). The “Setting” column indicates whether evidence is human- or machine-selected.
3 Assessing Explanation Transparency and Perceived Utility
We use Amazon Mechanical Turk to source workers to evaluate two qualities of the generated explanations: transparency, which assesses whether an explanation is citing the evidence passages appropriately; and utility, which measures the extent to which an explanation helps clarify the claim.
Transparency Evaluation
Assume we have a claim $c$, a veracity label $v$, $m$ evidence passages $E=\{e_{1},e_{2},e_{3},...,e_{m}\}$, and a generated explanation with $n$ sentences $X=\{x_{1},x_{2},x_{3},...,x_{n}\}$. [^3] Sentences are segmented with spaCy v3.7.2. A subset of sentences $X_{cit}=\{x_{i},x_{i+1},...,x_{j}\}\subseteq{X}$ contains inline citations (e.g. [6]), where ${1}\leq{i}\leq{j}\leq{n}$. We randomly select $e_{k}\in E$ (${1}\leq{k}\leq{m}$) and mask its inline citation marker in the explanation (e.g. you are wrong [6] $\rightarrow$ you are wrong), producing $X^{mask}$, the masked explanation. We denote $X^{mask}_{cit}\subseteq{X_{cit}}$ as the subset of explanation sentences without citation markers.
We next ask annotators to recover the masked sentences (see Appendix D for more annotation details). That is, annotators are presented with claim $c$, veracity $v$, full evidence set $E$, evidence passage $e_{k}$ and masked explanation ${X^{mask}}$, and they are asked to find all sentences that should cite $e_{k}$. In other words, the task is to recover $X_{cit}^{mask}$. Denoting their prediction as $X_{pred}^{mask}$, a perfect identification means $X_{pred}^{mask}=X_{cit}^{mask}$. Note that $X_{cit}^{mask}$ sometimes contains multiple sentences (e.g. when the explanation has 2 sentences that cite $e_{k}$) so this is not a trivial judgement. To measure the degree of overlap between $X_{pred}^{mask}$ and $X_{cit}^{mask}$, we compute set precision, recall, and F1 (see Appendix B).
Utility Evaluation
Utility measures to what extent the generated explanation clarifies a claim. Though a five-point Likert scale is commonly used (Liu et al., 2023; Gao et al., 2023), Ethayarajh and Jurafsky (2022) found that averaging using the Likert scale can result in a biased estimate. As such we use Direct Assessment (Graham et al., 2013) where annotators rate on a (continuous) sliding scale from 00100100100100, where $100$ = best. [^4] We ask annotators the following question: How helpful is the explanation in clarifying the truthfulness of the claim? We aggregate the judgements of multiple annotators for each explanation by computing the mean. [^5] We also applied a Bayesian model for utility score calibration, but it showed a similar tendency, so we decided to use the original scores. See Appendix F for details.
4 Experiments
Dataset
We use PolitiHop (Ostrowski et al., 2021) as our dataset for all experiments. An instance contains a claim, a veracity label, a set of evidence passages, and a subset of human-selected evidence passages. The human-selected evidence passages are a set of relevant passages that clarify the claim. For each claim, PolitiHop contains multiple human-selected evidence sets; we randomly select one for our experiments.
Large Language Models
@@ 4 Experiments
We experiment with three language models: GPT-4 OpenAI (2024) (gpt-4-0613, accessed Aug-Nov, 2023), GPT-3.5 Turbo (gpt-3.5-turbo-16k, accessed Aug-Nov, 2023) and LLaMA2-70B (Touvron et al., 2023). [^6] We also tested other models (LLaMA2-7B, FlanT5-xxl (Chung et al., 2024), Falcon-30B (Almazrouei et al., 2023) and MPT (Team, 2023)) but excluded them because these models generated repeated content with fabricated citations. Note that we always use the same LLM for both evidence selection and explanation generation, and as such the machine-selected evidence can also be interpreted as self-selected evidence. We generate explanations for 100 claims sampled from PolitiHop. See the Table 4 in the Appendix for generation statistics.
Evaluation
In addition to citation precision/recall/F1 for transparency and mean utility scores (Appendix B), we also compute the entropy of the distribution of human judgements for transparency. [^7] Concretely, an outcome corresponds to an explanation sentence, and the weight for an outcome (sentence) is the fraction of annotators who selected it. Entropy tells us how well the annotators agree with each other: a low entropy means annotators selected a similar set of sentences.
5 Results and Discussion
Though competitive, there’s still room for LLMs to improve on transparency.
As shown in Table 1, GPT-4 is the most transparent model (in terms of precision, recall, F1, and Entropy) by a comfortable margin, followed by GPT-35 and LLaMA2. Manual analysis reveals that LLaMA2 explanations often contain incorrect citations, such as [1][2][3] which are directly copied from the prompt/instruction, as well as random links. That said, even the best model (GPT-4 with machine-selected evidence) only produces an F1 of approximately 0.74, which means a good proportion of citations are still not faithful. There’s also substantial variance ($\pm$0.31), indicating the performance is far worse in the worst case scenario, limiting the real-world applicability.
Machine-selected evidences result in similar or better transparency and utility scores for explanations.
Table 1 shows that for GPT-4, machine-selected evidence produces substantially better explanations in terms of transparency (F1 = 0.74 vs. 0.63) and utility (76.34 vs. 66.86). For GPT-35 and LLaMA2-70b, the results are similar, [^8] The exception here is Entropy, where we see a much lower entropy for machine-selected evidence (0.21 vs. 0.41 for GPT-35; 0.21 vs. 0.34 for LLaMA2-70b); this result implies machine-selected evidence is somewhat better (as it leads to higher annotator agreement for recovering the masked sentences). suggesting that machine-selected evidence is at least as good as that selected by humans.
Machine-selected evidence is more comprehensive.
Our previous finding suggests two possibilities: (1) LLMs are tolerant to noisy evidence for explanation generation; or (2) machine-selected evidence is higher-quality in the first place. Looking at the evidence retrieval performance (Table 2) where we measure how well the machine-selected evidence set overlaps with the human-selected evidence set, we see that GPT-4 performs best compared to other LLMs, but in absolute terms (F1 = 0.47) the machine-selected set is still substantially different to the human-selected set. To understand the quality of machine-selected evidence better, we manually examine the GPT-4-selected evidence and found that GPT-4 selected more evidence passages on average (5 vs. 3 per claim) and as such the evidence set is more comprehensive. We also notice that there is sometimes redundancy in GPT-4 selected evidence passages, but this isn’t necessarily a negative outcome as it can further reinforce a key point in the explanation.
High utility doesn’t necessarily imply high transparency.
Although we found a a general correlation between utility and transparency (Table 1), GPT-35 achieved the highest utility score when using human-selected evidence (72.47), even though its transparency is much lower compared to GPT-4 (0.52 vs. 0.63). This demonstrates that utility and citation represent two distinct qualities and has an important implication: an explanation that appears helpful can actually still be misleading.
Transparency is hindered by LLM’s parametric knowledge.
We manually analyse instances with the lowest F1 (transparency) and found that the majority of them (60%) include fabricated statements, likely generated using the LLM’s parametric knowledge and not the evidence passages.
Model Precision Recall F1
GPT-4 0.40$\pm$0.21 0.75$\pm$0.23 0.47$\pm$0.20
GPT-35 0.34$\pm$0.22 0.60$\pm$0.35 0.39$\pm$0.22
LLaMA2-70b 0.29$\pm$0.18 0.68$\pm$0.32 0.36$\pm$0.18
Table 2: Evidence retrieval performance.
6 Related Work
Automated fact-checking aims to classify the veracity of a claim (Guo et al., 2022; Russo et al., 2023). However, debunking by giving a classification label (e.g. False) is not persuasive and can induce a “backfire” effect where the erroneous belief is reinforced (Lewandowsky et al., 2012; Guo et al., 2022). This motivates the task of justification or explanation generation for fact-checking. Explanation generation in the literature is framed as either extractive or abstractive summarization over evidence (Atanasova et al., 2020; Kotonya and Toni, 2020; Xing et al., 2022). But extractive methods struggle to produce explanations with sufficient context and abstractive methods are prone to hallucination (Russo et al., 2023). To address hallucination, explanation generation models now incorporate citations to sources (Huang and Chang, 2023)Liu et al. (2023) introduced the evaluation of citation quality and used human judges to audit the verifiability of popular commercial generative engines.  Gao et al. (2023) developed automatic metrics along three dimensions: fluency, correctness, and citation quality, to assesss LLM for question answering.
7 Conclusion
We explore the impact of human vs. machine-selected evidence for generating fact-checking explanations with LLMs. Surprisingly, we found that machine-selected evidence is either on-par or superior to human-selected evidence, depending on the LLM. Further analyses reveal that machine-selected evidence is more comprehensive and tends to contain redundant evidence (which help reinforce a key point). Our results also show that an explanation that appear helpful can be misleading in terms of how it cites the sources, highlighting the need to assess utility and transparency separately. Lastly, even though the best LLM exhibits encouraging performance, there’s still much room for improvement for explanation generation.
Limitation
This work only evaluates the transparency of the generated explanations, i.e. how faithful they are based on the sources. We want to note that this transparency evaluation does not assess factuality, which checks whether an explanation is factually correct. Additionally, while we tried our best to craft the prompts, there remains the question of whether transparency can be further improved with more prompt engineering.
The PolitiHop dataset may contain out-of-date information, given it was developed a few years ago. As such, this may conflict with the LLM’s parametric knowledge for certain claims. We saw this in our analyses where LLMs ignore evidence passages and generate statements based on their intrinsic knowledge. Also, the claims and evidence passages in PolitiHop are generally short single sentences. In practice, claims and evidence passages are likely to be longer.
Ethics Statement
Appendix A PolitiHop Dataset Sample
Table 3 shows an example from PolitiHop. As illustrated in  Figure 1, LLMs produce “Generated Explanation” that cite “Evidences”.
Claim: Shands Hospital of Gainesville Florida has confirmed its first case of coronavirus.
passage 1: Anglerfish may have a reputation for being among the creepier-looking ocean-dwellers..
passage 8 : What’s more, the picture referenced in the Facebook post alleging that anglerfish are typically 7 feet is taken from the Australian Museum’s 2012 exhibit titled “Deep Oceans”.
passage 10 : When the exhibit opened in June 2012, The Sydney Morning Herald reported on how the exhibit2̆019s team had created an “oversized anglerfish”…
Veracity: False
Evidences: {8,10}
Generated Explanation: The claim that the typical anglerfish is seven feet long is false. The confusion originates from a picture published in a 2012 Facebook post, which was actually taken from the Australian Museum’s exhibit “Deep Oceans”[8]. When this exhibit was inaugurated in the same year, The Sydney Morning Herald revealed that the exhibit’s team had developed an “oversized anglerfish”, indicating that the size of the anglerfish in the picture is not representative of the typical size for the species [10].
Table 3: An example based on PolitiHop with claim $c$, veracity $v$, associated evidences $E$. $X$ is the generated explanation that cites $E$.
Appendix B Transparency Evaluation Details

### v2

# Evaluating Evidence Attribution in Generated Fact Checking Explanations

@@ Abstract
Automated fact-checking systems often struggle with trustworthiness, as their generated explanations can include hallucinations. In this work, we explore evidence attribution for fact-checking explanation generation. We introduce a novel evaluation protocol, citation masking and recovery, to assess attribution quality in generated explanations. We implement our protocol using both human annotators and automatic annotators, and find that LLM annotation correlates with human annotation, suggesting that attribution assessment can be automated. Finally, our experiments reveal that: (1) the best-performing LLMs still generate explanations with inaccurate attributions; and (2) human-curated evidence is essential for generating better explanations. Code and data are available here: https://github.com/ruixing76/Transparent-FCExp.
Evaluating Evidence Attribution in Generated Fact Checking Explanations
@@ 1 Introduction
Claim: In 2018, for every "share" that a photograph of a sick boy gets, Facebook will donate one dollar.
Evidence passage 9: Facebook does not make donations based on the number of shares or comments a particular post or photo may receive.
Evidence passage 10 : The post is an example of a long-running scam that is employed to generate user interaction by falsely claiming that shares or comments will help pay for a sick child’s care.
Evidence passage 11 : In 2015, the Better Business Bureau released a statement about such scams…
Veracity: False
Selected evidence: {9, 10, 11}
Generated Explanation: The claim that Facebook will donate a dollar for every "share" that a photograph of a sick boy receives is false. It is crucial to note … a particular post or photo may receive [9]. Such posts are typically part of a long-running scam … will financially contribute to the care of a sick child [10]. Though it might initially appear benign,…, which in turn empowers them to disseminate further deceptions [11].
Table 1: An example based on PolitiHop with claim $c$, veracity $v$, evidence passages $E$. $X$ is the generated explanation that cites $E$.
The dissemination of online misinformation poses a significant threat to society, with consequences ranging from seeding skepticism and discrediting science, to endangering public health and safety. To verify online claims at scale, automated fact-checking systems have been proposed to classify claims based on their truthfulness (Guo et al., 2022). However, debunking by simply calling a claim “false” can lead to a backfire effect, whereby the belief in a false claim is further reinforced rather than hindered (Lewandowsky et al., 2012).
@@
As such, there’s growing research on generating textual explanations to justify the outcomes of fact-checking systems using large language models (LLMs) in various contexts (Eldifrawi et al., 2024; Wiegreffe et al., 2022). For automated fact-checking, the typical input is a claim and a list of retrieved evidence passages, from which a subset of evidence passages is selected and fed to the LLM for explanation generation. Table 1 presents a typical example of the fact-checking task with a generated explanation, and Figure 1 illustrates the process of generating these explanations.
@@
Figure 1: Automated explanation generation in fact-checking. Given a claim, its veracity and a list of evidence passages, a subset of these passages is selected, either by humans or machines, and input into a large language model (LLM) along with the claim to generate the explanation.
A challenge in this task is that the explanations must provide accurate attribution of their key statements (Nakov et al., 2021). In journalistic fact-checking practice, journalists curate their sources and ensure the accuracy of every cited statement (Baker and Fairbank, 2022). Machine-generated explanations should this follow convention to be trustworthy. But even if generated explanations contain attribution, how do we assess whether the attribution is faithful to the source evidence? Current evaluation methods are primarily based on the Attribution to Identified Source (AIS) framework (Rashkin et al., 2023), whereby an important feature is “explicature”: the meaning of statement $s$ in linguistic context $c$. Generating explicatures typically involves either costly human annotation or a non-trivial process of decontextualization, meaning it is rarely used. Consequently, most automated evaluation methods directly use Natural Lanugage Inference (NLI) between $c$ and $s$ to quantify AIS (Gao et al., 2023; Liu et al., 2023; Li et al., 2024).
In this work, we ground the notion of explicature by incorporating contextual information into the evaluation process. We introduce a novel evaluation protocol – Citation Masking and Citation recovery – to assess evidence attribution in generated explanations. The key idea is to mask the evidence citation in the generated explanation and ask annotators to recover them. The recovery process requires annotators to incorporate the linguistic context of a candidate statement during evaluation. This results in zero or more candidate statements, reflecting the complex nature of real-world attribution, and we hence frame it as a multi-label task. We first ask annotators to evaluate the attribution of a statement within the context of the full explanation. We then automate our protocol using NLI methods and LLMs as annotators. Additionally, we explore the influence of evidence selection on attribution quality, comparing human-curated and machine-selected evidence.
To summarise, our contributions are: (1) we introduce a new evaluation protocol for evidence attribution in generated explanations via citation masking and recovery; (2) we automate our protocol via LLMs and compare it with NLI models, showing that LLMs align better with human annotators; and (3) in terms of evidence source, we explore the influence of human- and machine-selected evidence on explanation generation and find that human-selected evidence produces explanations that are more accurate in terms of evidence attribution.
Figure 2: Our human assessment protocol: citation masking and citation recovery. Given the generated explanation and a list of evidence passages, we randomly masked one sentence $e_{k}$ and mask its inline citation marker. Annotators are then required to perform citation recovery and predict the masked citation sentence.
2 Related Work
Automated fact-checking aims to classify the veracity of a claim (Guo et al., 2022; Russo et al., 2023). However, debunking solely by providing a classification label (e.g. False) is neither effective nor convincing, and can induce a “backfire” effect where the erroneous belief is reinforced (Lewandowsky et al., 2012; Guo et al., 2022). This motivates the task of justification or explanation generation for fact-checking. Explanation generation in the literature is framed as either extractive or abstractive summarization over evidence (Atanasova et al., 2020; Kotonya and Toni, 2020; Xing et al., 2022). However, extractive methods struggle to produce explanations with sufficient context, and abstractive methods are prone to hallucination (Russo et al., 2023).
To alleviate hallucination, most existing work incorporates citations in the generation and evaluates their attribution relative to sources (Liu et al., 2023; Gao et al., 2023; Yue et al., 2023; Huang and Chang, 2024; Li et al., 2024). Liu et al. (2023) introduced evaluation of citation quality, and used human judges to audit the verifiability of popular LLMs.  Gao et al. (2023) developed automatic metrics along three dimensions – fluency, correctness, and citation quality – to assess LLMs for question answering. Yue et al. (2023) explored prompting LLMs and fine-tuning smaller LMs for automatic attribution evaluation. Most existing work is based on the Attribution to Identified Source (AIS) framework (Rashkin et al., 2023), where an annotator is asked whether the statement $s$ in the explanation is fully supported by cited evidence $c$ (According to evidence $c$, $s$.). An important assumption in AIS is explicature, an expanded statement that is interpretable out of context. Generating explicatures is a non-trivial process, often requiring costly human annotation to decontextualize statements so they can be interpreted in isolation. Most work directly adopts NLI without considering explicature (Liu et al., 2023; Gao et al., 2023). In this work, we attempt to bridge the gap by introducing a new evaluation protocol to evaluate evidence attribution in fact checking explanations.
3 Human Evaluation of Evidence Attribution
3.1 Dataset
We use PolitiHop (Ostrowski et al., 2021) as our dataset for all experiments. Table 1 shows an example instance, comprised of a claim, a veracity label, a list of evidence passages that provide additional relevant information about the claim, and selected evidence – a subset of passages selected by human annotators that is most relevant to the claim and collectively clarifies the veracity of the claim. [^1] In PolitiHop, the evidence in each claim can sometimes have multiple subsets (e.g. {8,10} and {6, 7, 8}); for these cases we randomly select one subset for our experiments.
3.2 Automatic Explanation Generation
@@
PolitiHop does not provide an explanation for each claim, and as such we need to generate the explanations for our experiments. We developed an LLM-based explanation generation system which assumes the following as input: (1) a claim; (2) a list of evidence passages; and (3) the veracity of the claim (e.g. False). The rationale for assuming the veracity label as input is we see our explanation generation system being applied to a fact-checking classification system to provide a layer of interpretability in a real-world application. Because the set of evidence passages is typically large and not all of them are relevant, we propose a pipeline approach where we first do evidence selection and then feed the selected evidence, claim, and veracity label to an LLM to generate the explanation. Figure 1 illustrates the overall process.
@@
We experiment with two methods for evidence selection: (1) Human: we use the original human-selected evidence in Politihop; and (2) Machine: we (one-shot) prompt an LLM to select the most relevant evidence passages (see C14 for more details). Note that the evidence passages are numbered, which is important as the explanation cites them for key statements. We explore these two methods to understand the impact of this selection process – whether hand-curated evidence is necessary for explanation generation, or if machine-selected evidence is sufficient.
@@
Given a claim, veracity label, and selected evidence (machine- or human-selected), we (zero-shot) prompt an LLM to generate an explanation to clarify the claim. The prompt explicitly asks it to use in-line citations, making it clear that all given evidence passages should be used (see “Generated Explanation” in Figure 1; the full prompt is given in Appendix Table 12). We test a range of LLMs in our experiments, noting that we always use the same LLM for evidence selection and explanation generation. We generate explanations for 100 claims sampled from PolitiHop.
Evi Src Gen Claim Len Evi Size Exp Len
Human GPT-4 22.32 3.16 140.58
GPT-3.5
LLaMA2
Machine GPT-4 22.32 5.07 171.26
GPT-3.5 4.68 175.89
LLaMA2 6.24 214.02
Table 2: Statistics of the data and generated explanations. The “Evi Src” column indicates whether evidence is human- or machine-selected. “Gen” means the generator model. “Claim Len” and “Exp (Explanation) Len” refer to token length, which is tokenized by OpenAI’s tiktoken v0.7.0. Evi (Evidence) Size refers to the number of selected evidence passages.
Models
@@
For the evidence selector and explanation generator, we experiment with three language models: GPT-4 (OpenAI, 2024), GPT-3.5, and LLaMA2 (Touvron et al., 2023). [^2] Detailed model version: GPT-4 = gpt-4-0613, GPT-3.5 = gpt-3.5-turbo, and LLaMA2 = Llama-2-7b-chat. Note that we always use the same LLM for both evidence selection and explanation generation, and as such the machine-selected evidence can also be interpreted as self-selected evidence. Table 2 presents data statistics for the explanation generation process. In general, we can see that the machine-selected evidence set tends to be larger (i.e. it includes more passages), and the explanations generated using machine-selected evidence are also longer.
3.3 Human Assessment Protocol
In this section, we introduce citation masking and recovery, a new approach for assessing the quality of evidence attribution.
Citation Masking
Assume we have a claim $c$, a veracity label $v$, $m$ evidence passages $E=\{e_{1},e_{2},e_{3},...,e_{m}\}$, and a generated explanation with $n$ sentences $X=\{x_{1},x_{2},x_{3},...,x_{n}\}$. [^3] Sentences are segmented with spaCy v3.7.2.
A subset of sentences $X_{cit}=\{x_{i},x_{i+1},...,x_{j}\}\subseteq{X}$ contains inline citations (e.g. [6]), [^4] Technically there is a possibility that $X=\emptyset$ when the LLM fails to cite any evidence passages, but we did not see this in practice. where ${1}\leq{i}\leq{j}\leq{n}$. We randomly select $e_{k}\in E$ (${1}\leq{k}\leq{m}$) and mask its inline citation marker in the explanation (e.g. you are wrong [6] $\rightarrow$ you are wrong), producing $X^{mask}$, the masked explanation. We denote $X^{mask}_{cit}\subseteq{X_{cit}}$ as the subset of explanation sentences where their citation markers are removed.
Citation Recovery
We next ask annotators to recover the masked sentences. That is, annotators are presented with claim $c$, veracity $v$, full evidence set $E$, evidence passage $e_{k}$, and masked explanation ${X^{mask}}$, and they are asked to find all sentences that should cite $e_{k}$. In other words, the task is to recover $X_{cit}^{mask}$. Denoting their prediction as $X_{pred}^{mask}$, a perfect identification means $X_{pred}^{mask}=X_{cit}^{mask}$. Note that $X_{cit}^{mask}$ sometimes contains multiple sentences (e.g. when the explanation has 2 sentences that cite $e_{k}$) so this is a multi-label classification problem. To measure the degree of overlap between $X_{pred}^{mask}$ and $X_{cit}^{mask}$, we compute set precision, recall, and F1. A high F1 performance indicates the explanation is accurate in its attribution and faithful to the source evidence.
Attribution score computation